OpenStudy (anonymous):

Please, if someone could please help me with this problem... If np ≥ 5 and nq ≥ 5, estimate P (more than 4) with n = 11 and p = 0.7 by using the normal distribution as an approximation to the binomial distribution; if np < 5 or nq < 5, then state that the normal approximation is not suitable. a. P(more than 4) = b. The normal distribution cannot be used.

OpenStudy (amistre64):

well, what is np and nq?

OpenStudy (anonymous):

np= 11*0.7 nq=11*4

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

p= .7 q = 1-p = .3 11(.7) and 11(.3)

OpenStudy (anonymous):

7.7 and 3.3

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

is 3.3 >= 5 ?